# Yuchen Zhang

### v2

@@ 6 Conclusion
This study discloses two critical limitations in current multimedia manipulation detection: underestimation of dynamic semantic deception risks posed by MLLMs and the unrealistic, semantically incoherent misalignment artifacts among existing benchmarks. To address these challenges, we construct the MLLM-Driven Synthetic Multimodal (MDSM) dataset and the Artifact-aware Manipulation Diagnosis (AMD) framework to address this new and challenging problem. AMD integrates Artifact Pre-perception Encoding and Manipulation-Oriented Reasoning to enhance detection of MLLM-generated multimodal disinformation. Comprehensive experiments demonstrate the framework’s superior generalization capabilities, validating its effectiveness as a unified solution for combating advanced MLLM-driven deception.
Appendix
Appendix Contents
Appendix A Assessing Public Risks of MLLM Narratives - A Human Evaluation
To assess human ability to identify multimodal misinformation generated by MLLMs in combination with image manipulation models, we design a human evaluation study based on 100 image-text pairs sampled from our MDSM dataset. As illustrated in Fig. 5, each test sample belonged to one of four categories: both image and text are original (Orig), only image is manipulated (Fake Image, F-I), only text is manipulated (Fake Text, F-T), and both image and text are manipulated (Fake Image & Text, F-I&T).
We recruit 15 adult volunteers, all holding at least a bachelor’s degree, to participate in the evaluation. As summarized in Fig. 6a, the accuracy of identifying Orig, F-I, F-T, and F-I&T samples was 53.19%, 13.22%, 20.18%, and 22.42%, respectively. Notably, all manipulated categories exhibit low recognition accuracy, with none exceeding 23%. To better visualize classification performance and patterns of confusion, we construct a row-normalized confusion matrix (Fig. 6b). The matrix reveals that among all manipulated types, F-I&T samples were most frequently misclassified as original real news, with a false-negative rate of 46.43%. This finding suggests that fake news jointly generated by MLLMs and image editing models can achieve high semantic coherence and pose a significant threat in terms of deception.
Considering real-world scenarios, however, humans often do not need to identify which modality is manipulated; it is often sufficient to detect the presence of any form of misinformation to avoid being misled. Therefore, we introduce two binary metrics to quantify this ability: Overall Recall and False Alarm Rate. The Overall Recall is defined as the proportion of manipulated items (F-I, F-T, F-I&T) correctly identified as any type of fake (i.e., not labeled as Orig). The False Alarm Rate refers to the proportion of original items (Orig) incorrectly identified as any form of fake. We further compute the per-category fake detection rate for F-I, F-T, and F-I&T individually. The results are illustrated in Fig. 6c. The analysis shows that participants identify 56.7% of the manipulated samples as fake under this relaxed criterion. However, they also falsely flag 48.8% of the original news items as fake. This elevated false alarm rate indicates a conservative judgment tendency in the testing environment, implying that the actual detection rate in real-world conditions may be significantly lower than 56.7%. Among manipulated categories, F-I samples has the highest detection rate at 64.5%, compared to 54.5% for F-T and 53.6 for F-I&T. This discrepancy suggests that MLLM-generated textual fabrications in our MDSM dataset are particularly deceptive and challenging to identify.
Following the same evaluation setup, we also conducted an investigation on the $\text{DGM}^{4}$ dataset. The results, shown in Fig. 6d-f, indicate that the $\text{DGM}^{4}$ dataset exhibits similarly strong deception. The recognition accuracy for all manipulation samples does not exceed 36% (Fig. 6d). For the samples involving text manipulation in $\text{DGM}^{4}$, the recognition accuracies for F-I and F-I&T are 28.25% and 35.40%, respectively, both higher than the 20.18% and 22.42% for MDSM. This suggests that the fabricated text generated by MLLMs, as considered in MDSM, is more likely to mislead the general public.
In summary, the findings highlight a generally low human sensitivity to misinformation generated by MLLMs and image editing systems, especially in cases where multimodal manipulations are semantically consistent. This underscores the real-world threat posed by MLLM-involved fake news and points to the urgent need for robust automatic misinformation detection systems to mitigate societal harm and support informed decision-making.
Figure 5: The user interface of the human evaluation study where each participant is given pairs of news images and caption and asked to determine whether they are manipulated or not.
Figure 6: Human evaluation statistics on multimodal fake news identification. (a) Per-class accuracy across four types of image-text pairs on MDSM. (b) Row-normalized confusion matrix showing classification tendencies on MDSM. (c) Human perception of manipulated news and misclassification of real content on MDSM. (d) Per-class accuracy across four types of image-text pairs on $\text{DGM}^{4}$. (e) Row-normalized confusion matrix showing classification tendencies on $\text{DGM}^{4}$. (f) Human perception of manipulated news and misclassification of real content on $\text{DGM}^{4}$.
Appendix B Distribution of MLLM-generated Texts
To evaluate the quality of fake corpus generated by MLLM, we compare their textual distributions against the authentic news corpus in MDSM. We conduct a statistical analysis using SpaCy [explosion_spacy] and TextBlob [sloria_textblob] across five linguistic dimensions: (1) average sentence length (syntactic complexity), (2) top-10 frequent words (topical and lexical overlap), (3) noun–verb ratio (stylistic tendencies), (4) type–token ratio (lexical diversity), and (5) average sentiment score (tonal neutrality).
As shown in Tab. 5, the distributions of generated texts exhibit strong proximity to real news. The average sentence length and noun–verb ratio are nearly identical, indicating that the MLLM outputs capture comparable syntactic rhythm and stylistic balance. Frequent word distributions overlap substantially (e.g., ’new’, ’president’, ’mr’, ’one’, ’join’), as further illustrated by the word clouds in Fig. 7, reflecting clear topical alignment. The type–token ratio of generated texts is 0.13, only slightly lower than authentic news’s 0.15, which is expected given structured prompts and reduced lexical randomness. Sentiment scores in both corpora remain close to neutral and slightly positive (the authentic is 0.04 and the MLLM-genarated is 0.08), consistent with the stylistic norms of mainstream journalism.
Overall, these results confirm that MLLMs, when guided by carefully designed instructions, can produce texts that closely mimic the linguistic distribution of authentic news across multiple dimensions.
Table 5: Statistical comparison results between real and MLLM-generated text.
Metric Authentic Corpus MLLM-Generated Corpus
Average Sentence Length 20.79 21.98
Top 10 Frequent Words left, said, new, president,mr, last, center, one, right, join new, join, us, event, president, one, seen, hosting, mr, york
Noun–Verb Ratio 2.20 2.00
Type–Token Ratio 0.15 0.13
Average Sentiment Score 0.04 0.08
Figure 7: Word Clouds of Authentic and MLLM-Generated Corpus.
Appendix C Discussion of MDSM and $\text{DGM}^{4}$
C.1 Contribution of MDSM
The core contribution of our MDSM dataset is threefold: (1) It addresses the under-explored yet critical threat of misinformation crafted by MLLMs; (2) It provides the first dedicated benchmark for detecting semantically-aligned news manipulation – a significantly more realistic and challenging scenario; (3) MDSM enhances richness and scale (almost 2 times larger than $\text{DGM}^{4}$) by collecting broader media types, news topics, and diverse sources (5 domains across 3 sources: NYT, BBC, USA Today, Guardian, Wash.) compared to prior datasets like $\text{DGM}^{4}$ (4 domains, 1 source).
Crucially, MLLM-drafted narratives are highly deceptive, as evidenced by human evaluators achieving only 22.42% accuracy (Fig. 6). This starkly contrasts with the misaligned image-text contexts in datasets like $\text{DGM}^{4}$, where human detection is far easier (35.40% accuracy). This gap underscores the unique challenge MDSM addresses.
Therefore, MDSM establishes a realistic benchmark specifically designed for the emerging threat of MLLM-generated disinformation, driving essential progress in cross-modal forgery detection. With the boom of MLLMs, we need to consider the threat of their malicious use in social multimodal news to deceive the public. From this aspect, our MDSM is a timely contribution to promote this direction of research.
C.2 Challenge Level of MDSM and $\text{DGM}^{4}$
Table 6: Discussion experiments on MDSM and $\text{DGM}^{4}$
Len. $\text{MDSM}_{\text{Gs}}$ $\text{DGM}^{4}_{\text{G}}$
ACC mAP mIoU ACC mAP mIoU
NYT 72.26 40.28 44.70 $\backslash \\backslash$
Guardian 87.07 78.32 80.26 84.61 68.50 81.24
USA 73.32 40.18 43.07 70.62 43.20 41.99
Wash. 70.16 41.17 42.17 70.28 43.36 39.05
BBC 74.63 50.64 46.93 72.37 56.57 45.20
AVG 75.49 50.12 51.43 74.47 52.91 51.87
(a) Comparation on $\text{MDSM}_{\text{Gs}}$ and $\text{DGM}^{4}_{\text{G}}$.
Setting HAMMER AMD
ACC mAP mIoU ACC mAP mIoU
$\text{DGM}^{4}$ 93.06 85.19 75.86 90.23 83.17 76.19
$\text{MDSM}_{r}$ 91.27 62.79 61.24 91.83 81.96 78.63
Performance Gap -1.79 -22.40 -14.62 +1.60 -1.21 +2.44
(b) HAMMER and AMD’s performance. Performance Gap is from $\text{DGM}^{4}$ to $\text{MDSM}_{r}$
Methods Test Dataset $\text{AVG}_{\text{ACC}}$
MDSM COSMOS [COSMOS_AAAI] MiRAGe [huang-etal-2024-miragenews]
ACC mAP mIoU ACC ACC
ViLT 39.62 21.18 21.93 51.27 31.18 40.69
HAMMER 46.02 24.88 35.19 57.14 34.12 45.76
HAMMER++ 46.17 24.78 33.78 57.79 34.56 46.17
FKA-Owl 54.23 32.48 36.76 57.02 38.76 50.00
AMD (Ours) 54.75 31.49 43.68 57.16 39.68 50.53
(c) Zero-Shot transfer performance of $\text{DGM}^{4}$-trained models on other benchmarks. $\text{AVG}_{\text{ACC}}$ is the average accuracy across all datasets in each row.
It is important to clarify that the motivation behind constructing MDSM is not to create a dataset more challenging than $\text{DGM}^{4}$. Rather, MDSM addresses a fundamentally distinct threat: MLLM-crafted, semantically aligned multimodal disinformation, which remains an underexplored vulnerability in existing benchmarks. This represents a critical and emerging risk paradigm that necessitates dedicated investigation.
To enable a fair comparison, we reduce the MDSM-Guardian subset ($\text{MDSM}_{\text{Gs}}$) to match the 103k samples of the $\text{DGM}^{4}$-Guardian subset ($\text{DGM}^{4}_{\text{G}}$). Next, we train AMD on this subset and present the results in Tab. 6a. The $\text{MDSM}_{\text{Gs}}$-trained AMD achieves 75.49 ACC / 50.12 mAP / 51.43 mIoU, which are comparable to the results on $\text{DGM}^{4}$ (74.47 / 52.91 / 51.87, respectively). These comparable results suggest that MDSM and $\text{DGM}^{4}$ present similar levels of challenge under controlled conditions.
Additionally, we reduce the size of MDSM to match that of $\text{DGM}^{4}$, denoting this subset as $\text{MDSM}_{\text{r}}$, and conduct an evaluation without domain shift. As shown in Tab. 6b, AMD demonstrates strong performance on both datasets, with the performance gap within $\pm 3$. However, HAMMER exhibits substantial performance degradation on $\text{MDSM}_{\text{r}}$, particularly in terms of mAP (22.40 points drop) and mIoU (14.62 points drop). These results indicate that existing frameworks struggle to effectively handle semantically aligned manipulations.
The above experimental results indicate that MDSM and $\text{DGM}^{4}$ have similar levels of challenge, and MDSM is slightly more difficult in distinguishing manipulation types at a finer granularity. Additionally, our human verification results confirm this (the human discrimination accuracy rates for MDSM and $\text{DGM}^{4}$ are 22.42% and 35.69%, respectively).
C.3 Cross Evaluation among MDSM and $\text{DGM}^{4}$
The evaluation of our MDSM-trained models on $\text{DGM}^{4}$ and other benchmarks is reported in Tab. 10, where AMD achieves the best overall performance. We further train the models on $\text{DGM}^{4}$ and test them on MDSM and other benchmarks. As shown in Tab. 6c, AMD consistently demonstrates strong generalization to unseen cases, achieving the highest performance with an MDSM-ACC of 54.75, MDSM-mIoU of 43.68, and MiRAGE-ACC of 39.68, while also attaining the best average ACC performance of 50.53. Notably, on the MiRAGE dataset, all MDSM-trained models (except ViLT) achieve an accuracy above 50 (Tab. 10), whereas $\text{DGM}^{4}$-trained models consistently fall below 40. These results highlight the significant potential of MDSM-trained models in detecting purely generated multimodal disinformation.
Appendix D Prompt Paradigm
D.1 Prompt for AMD
@@
D.2 Prompt for General-purpose Model
To ensure fairer testing and more credible results for general-purpose models in Tab.2, we enhanced the invocation of general-purpose models by adding more detailed descriptions to the AMD prompt, as follows:
@@ D.2 Prompt for General-purpose Model
Appendix E Experimental Setup
E.1 Implementation Details
All experiments are conducted on 4 NVIDIA GeForce RTX 4090 GPUs using Distributed Data Parallel (DDP) training in PyTorch. The image encoder $\mathcal{E}_{v}$ is based on DaViT [ding2022davit], with Florence-2-B [Florence2_CVPR] serving as the backbone. The APE $\mathcal{E}_{m}^{p}$ is based on the Florence-2 encoder and remains frozen during training. The classifiers and bounding box (bbox) detector consist of two Multi-Layer Perceptron layers, with output dimensions of 2 and 4, respectively. For manipulation detection guidance, the Dual-Branch Manipulation shares a common classifier.
The training images are resized to $224\times 224$ and undergo random horizontal flipping. The batch size per GPU (per_GPU_bs) is set to 6, and the model is trained for 12 epochs. We use the AdamW optimizer [AdamW_ICLR] with an initial learning rate of $1e^{-7}$ and a weight decay of 0.01. A cosine learning rate scheduler with a warm-up phase is applied, gradually increasing the learning rate to $1e^{-6}$ in the first 1000 steps, and then decaying it to $1e^{-7}$ throughout training. Our code will be released to provide further implementation details.
E.2 Baselines
@@
  • HAMMER  [DGM4] is a pioneering model for the multi-modal manipulation detection and grounding. It employs two unimodal encoders to extract visual and textual forgery features, which are then aligned through contrastive learning. Following this, a multi-branch transformer architecture with two specialized decoders is utilized for manipulation detection and grounding.
  • HAMMER++  [DGM4_TPAMI] is a more powerful model that builds upon HAMMER by integrating contrastive learning from both global and local perspectives.
  • FKA-Owl  [liu2024fkaowl] is another pioneering model designed for large vision-language models to perform multi-modal fake news detection, and it demonstrates outstanding cross-domain performance. Since FKA-Owl does not support fine-grained classification tasks, we fine-tuned it using the same prompts as those used for AMD.
  • ViLT  [ViLT_ICML], for the multi-modal learning approach, is a representative single-stream method where cross-modal interaction layers operate on the concatenation of image and text inputs. For adaptation, We add classification and detection heads to the corresponding outputs of the model.
E.3 Evaluation Metrics
To comprehensively evaluate our proposed MDSM, we follow the rigorous evaluation protocols and metrics outlined in [DGM4] for all manipulation detection and grounding tasks. The detailed evaluation setup is organized as follows:
@@ E.3 Evaluation Metrics
  • Multi-Label Classification. For multi-label classification tasks, we employ mean Average Precision (mAP). This metric effectively captures the average performance across all labels, providing a comprehensive assessment of multi-dimensional manipulation type classification accuracy.

### v4

@@ 5 Conclusion
This study discloses two critical limitations in current multimedia manipulation detection: underestimation of dynamic semantic deception risks posed by MLLMs and the unrealistic, semantically incoherent misalignment artifacts among existing benchmarks. To address these challenges, we construct the MLLM-Driven Synthetic Multimodal (MDSM) dataset and the Artifact-aware Manipulation Diagnosis (AMD) framework to address this new and challenging problem. MDSM contains over 441k semantically aligned image-text pairs from five major media domains, providing a solid foundation for training robust misinformation detection models. AMD integrates Artifact Pre-perception Encoding and Manipulation-Oriented Reasoning to enhance detection of MLLM-generated multimodal disinformation. Comprehensive Experiments demonstrate the framework’s strong generalization and effectiveness against MLLM-driven deception.
Acknowledgment
This work was supported by the National Natural Science Foundation of China (NSFC) under Grant No. 62302140, and the National Key Research and Development Program of China under Grant No. 2023YFC3321600. The authors also gratefully acknowledge the support from the Guangdong Basic and Applied Basic Research Foundation (2025A1515012281), the Nanjing Municipal Science and Technology Bureau (202401035), and the University of Macau (MYRG-GRG2024-00077-FST-UMDF).
Supplementary Material
6 Related Work
Deepfake Detection. The rapid progress of generative models and the surge in synthetic content have accelerated advances in Deepfake detection. Existing work spans unimodal and multimodal approaches. Unimodal, including image-based [60, 21] and text-based [44, 14] approaches, already achieve strong results. With the rise of Multimodal Large Language Models (MLLMs), multimodal Deepfake detection has received increasing attention [24, 40, 23]. Regarding datasets, Shao et al. [40] introduced the pioneering DGM4 benchmark for multimodal manipulation detection and grounding. However, its manipulations are rule-based, leading to semantically fragmented image-text discrepancies that do not accurately reflect real-world misinformation. MMFakeBench [24] recognized this limitation and proposed generating semantically aligned news images using text-to-image models. Yet such semantically matched samples constitute only 30% of its fake subset, and the dataset contains merely 11k samples, limiting its utility for training robust detectors. Existing Deepfake datasets also fail to consider the risk of semantically coherent but misleading text generated by modern MLLMs. On the methodological side, HAMMER [40] integrates contrastive learning to build a detector capable of classifying manipulation types and grounding manipulated regions, but it does not address cross-domain robustness. Beyond conventional multimodal detectors, FKA-Owl [23] employs a 7B-scale MLLM with several architectural modifications to enhance generalization. However, it is trained on DGM4, where text manipulations follow fixed editing rules rather than being synthesized by MLLMs, making it unsuitable for detecting more subtle, semantically aligned misinformation produced by modern models. Moreover, FKA-Owl performs only binary real/fake classification without fine-grained manipulation type prediction or grounding, and its large backbone and heavy architectural design result in substantially increased computational cost and slower inference.
Multi-Modal Large Language Model. In recent years, Multi-Modal Large Language Models have emerged as a crucial technology for understanding and reasoning across multiple modalities, particularly text and images. By extending the capabilities of Large Language Models (LLMs) to incorporate visual inputs, these models have demonstrated outstanding performance in tasks such as image captioning and visual question answering. CLIP  [35] and ALIGN  [17] leveraged contrastive learning to align visual and textual representations, enabling efficient zero-shot vision-language understanding. Subsequently, models such as Flamingo  [2] and BLIP-2  [20] have introduced vision-language transformers, integrating pre-trained LLMs with vision encoders to enhance cross-modal reasoning and generative capabilities. More recently, GPT-4V  [30] and Florence-2  [5] have significantly enhanced the potential of MLLMs in tackling complex multi-modal tasks by leveraging a more efficient framework and larger-scale pre-training data. A key advantage of MLLMs is their acquisition of extensive world knowledge through large-scale pretraining, which substantially strengthens their reasoning abilities in downstream tasks. Such knowledge not only enhances cross-modal understanding but also proves essential for handling challenging problems, including misinformation detection.
7 Prompt Paradigm
7.1 Prompt for AMD
@@
7.2 Prompt for General-purpose Model
To ensure fairer testing and more credible results for general-purpose models (Tab.2 in the main paper), we enhanced the invocation of general-purpose models by adding more detailed descriptions to the AMD prompt, as follows:
@@ 7.2 Prompt for General-purpose Model
8 Experimental Setup
8.1 Implementation Details
All experiments are conducted on 4 NVIDIA GeForce RTX 4090 GPUs using Distributed Data Parallel (DDP) training in PyTorch. The image encoder $\mathcal{E}_{v}$ is based on DaViT [11], with Florence-2-B [5] serving as the backbone. The APE $\mathcal{E}_{m}^{p}$ is based on the Florence-2 encoder and remains frozen during training. Thus, in the APE stage, only the artifact token, classifier head, and attention pooling module are jointly trained. The classifiers and bounding box (bbox) detector consist of two Multi-Layer Perceptron layers, with output dimensions of 2 and 4, respectively. For manipulation detection guidance, the Dual-Branch Manipulation shares a common classifier.
The training images are resized to $224\times 224$ and undergo random horizontal flipping. The batch size per GPU is set to 6, and the model is trained for 12 epochs. We use the AdamW optimizer [26] with an initial learning rate of $1e^{-7}$ and a weight decay of 0.01. A cosine learning rate scheduler with a warm-up phase is applied, gradually increasing the learning rate to $1e^{-6}$ in the first 1000 steps, and then decaying it to $1e^{-7}$ throughout training. Our code will be released to provide further implementation details.
8.2 Baselines
@@
  • HAMMER  [40] is a pioneering model for the multi-modal manipulation detection and grounding. It employs two unimodal encoders to extract visual and textual forgery features, which are then aligned through contrastive learning. Following this, a multi-branch transformer architecture with two specialized decoders is utilized for manipulation detection and grounding.
  • HAMMER++  [41] is a more powerful model that builds upon HAMMER by integrating contrastive learning from both global and local perspectives.
  • FKA-Owl  [23] is another pioneering model designed for large vision-language models to perform multi-modal fake news detection, and it demonstrates outstanding cross-domain performance. Since FKA-Owl does not support fine-grained classification tasks, we fine-tuned it using the same prompts as those used for AMD.
  • ViLT  [18], for the multi-modal learning approach, is a representative single-stream method where cross-modal interaction layers operate on the concatenation of image and text inputs. For adaptation, We add classification and detection heads to the corresponding outputs of the model.
8.3 Evaluation Metrics
To comprehensively evaluate our proposed MDSM, we follow the rigorous evaluation protocols and metrics outlined in [40] for all manipulation detection and grounding tasks. The detailed evaluation setup is organized as follows:
@@ 8.3 Evaluation Metrics
  • Multi-Label Classification. For multi-label classification tasks, we employ the mean Average Precision (mAP), which measures the per-class average precision and then takes the arithmetic mean across all manipulation types. This macro-averaged mAP provides a comprehensive evaluation of the model’s overall performance across different manipulation types.